# Dual thermal pseudo-critical features in a spin-1/2 Ising chain with twin-diamond geometry

### v2

@@ IV.1 Entropy and specific heat
In Fig. 4 is reported the low-temperature entropy landscape in the $(\mu_{{\rm B}}g_{0}B/J_{1},\,k_{{\rm B}}T/J_{1})$ plane for $g_{1}\neq g_{0}$. Panel (a), obtained for $J_{0}/J_{1}=-0.01$, $J_{2}/J_{1}=1.2$ and $g_{1}/g_{0}=0.7$, shows three distinct low-temperature domains that reproduce the structure of the ground-state phase diagram. A narrow green plateau associated with the $\mathrm{FR}_{1}$ phase appears between the FI and $\mathrm{FR}_{2}$ sectors, and its boundary acquires a negative slope due to the asymmetric Zeeman response. In this region the entropy ridges separating the ordered and frustrated sectors mark the emergence of two pseudo-transition scales, one related to the competition between FI and $\mathrm{FR}_{1}$ and another associated with the competition between $\mathrm{FR}_{1}$ and $\mathrm{FR}_{2}$. These features become progressively smoother as the temperature increases. Panel (b), corresponding to $J_{0}/J_{1}=-0.01$, $J_{2}/J_{1}=2.4$ and $g_{1}/g_{0}=1.3$, shows the opposite situation where the frustrated boundaries tilt with positive slope, reflecting the opposite imbalance in the magnetic response. The pseudo-transition ridges remain visible, indicating that the entropic signatures of the frustrated manifold are robust against variations in the ratio $g_{1}/g_{0}$.
@@
Figure 4: Density plot of the entropy in the $(\mu_{{\rm B}}g_{0}B/J_{1},\,k_{{\rm B}}T/J_{1})$ plane, for a weak ferromagnetic coupling $J_{0}$. (a) $J_{0}/J_{1}=-0.01$, $J_{2}/J_{1}=1.2$ and $g_{1}/g_{0}=0.7$; (b) Similarly for $J_{0}/J_{1}=-0.01$, $J_{2}/J_{1}=2.4$ and $g_{1}/g_{0}=1.3$.
@@
Figure 5 shows the temperature dependence of the entropy and specific heat for two representative parameter sets reported in the caption of Fig.5, illustrating the emergence of two distinct pseudo-critical temperatures. These scales are associated with the $\mathrm{FI}$ to $\mathrm{FR}_{1}$ crossover, with pseudo-critical temperature given by Eq.(33) at $k_{{\rm B}}T_{p_{1}}/J_{1}\approx 0.0504764$, and with the $\mathrm{FR}_{1}$ to $\mathrm{FR}_{2}$ crossover, with pseudo-critical temperature given by Eq. (34) at $k_{{\rm B}}T_{p_{2}}/J_{1}\approx 0.0788948$. Panels (a) and (b) display the entropy curves, which show two smooth but clearly separated increases as temperature rises. The lower increase at $T_{p_{1}}$ corresponds to the partial activation of the $\mathrm{FR}_{1}$ manifold, whereas the higher increase at $T_{p_{2}}$ signals access to the fully frustrated $\mathrm{FR}_{2}$ sector.
@@
Panels (c) and (d) present the corresponding specific heat $C(T)$, which develops two continuous maxima at $T_{p_{1}}$ and $T_{p_{2}}$. The logarithmic insets make the locations of these maxima particularly visible. The first maximum marks the crossover between the ordered FI region and the partially frustrated $\mathrm{FR}_{1}$ configurations, while the second reflects the competition between $\mathrm{FR}_{1}$ and $\mathrm{FR}_{2}$. We can observe that the peak at $T_{p_{1}}$ is higher and sharper than the second peak at $T_{p_{2}}$. This behavior is consistent with the pseudo-critical features reported in the literature(sergio18, ; Rojas2019, ; katarina, ; verissimo, ).
@@
The clear separation between these double thermal anomalies shows that the model supports two well defined pseudo-critical temperature scales, each associated with a distinct degeneracy change in the ground-state structure.
@@ IV.2 Magnetic properties
Panels (a) and (b) display $\langle s_{k}\rangle$, $\langle S_{k}\rangle$ and $M$ as functions of temperature. All three curves show two smooth but clearly separated steps. The first step, located at the pseudo-critical temperature $T_{p_{1}}$, reflects the partial activation of the $\mathrm{FR}_{1}$ configurations, which introduce local frustration while keeping the nodal subsystem partially aligned. The second step, located at $T_{p_{2}}$, marks the activation of the fully frustrated $\mathrm{FR}_{2}$ manifold, where each dimer contributes an additional independent degree of freedom.
@@
Panels (c) and (d) of Fig. 6 show $\chi_{0}$, $\chi_{1}$, and $\chi_{t}$ on a logarithmic scale. All three susceptibilities develop two pronounced peaks at $T_{p_{1}}$ and $T_{p_{2}}$. These peaks coincide with the pseudo-critical temperatures identified from the entropy and specific-heat analysis. Although the peaks appear sharp, they remain finite and represent the typical thermal signatures of pseudo-transitions in frustrated one-dimensional systems. Again, the anomalies occur at the same temperatures identified from the entropy and specific heat, and the peak at $T_{p_{1}}$ is higher and sharper than the one at $T_{p_{2}}$. This confirms the consistency of the dual pseudo-transition behavior in the CTDC model.
@@ IV.3 Eigenvalues properties and correlation length
Figure 7(a) shows the temperature dependence of the relative eigenvalues $|\lambda_{j}|/\lambda_{1}$ for the parameter set $J_{2}/J_{1}=1.24$, $J_{0}/J_{1}=-0.01$, $\mu_{\mathrm{B}}g_{0}B/J_{1}=1.1$, and $g_{1}/g_{0}=0.7$. Although the eigenvalues satisfy the ordering $\lambda_{1}>\lambda_{2}>\lambda_{3}>\lambda_{4}$, the subleading eigenvalue in modulus is not $\lambda_{2}$ but rather $|\lambda_{4}|$ (see appendix A). This behavior is illustrated by the red curve $|\lambda_{4}|/\lambda_{1}$, while the green and blue curves correspond to $\lambda_{2}/\lambda_{1}$ and $\lambda_{3}/\lambda_{1}$, respectively. The ratio $|\lambda_{4}|/\lambda_{1}$ exhibits a broad plateau between the two pseudo-critical temperatures, indicating that the same low-energy sector remains dominant throughout this interval. In contrast, the ratio $\lambda_{2}/\lambda_{1}$ displays two pronounced extrema located precisely at the pseudo-critical temperatures $T_{p_{1}}$ and $T_{p_{2}}$, reflecting the successive competition between distinct thermodynamic sectors.
For the same parameters, Fig. 7(b) shows the quantities $\xi_{j}=\bigl[\ln(\lambda_{1}/|\lambda_{j}|)\bigr]^{-1}$ as functions of temperature on a logarithmic scale. Since $|\lambda_{4}|$ is the largest subleading eigenvalue in modulus, the spectral correlation length is given by $\xi=\xi_{4}=\bigl[\ln(\lambda_{1}/|\lambda_{4}|)\bigr]^{-1}$. However, this quantity does not exhibit signatures of the pseudo-critical behavior. By contrast, the quantity $\xi_{2}=\bigl[\ln(\lambda_{1}/\lambda_{2})\bigr]^{-1}$, although not a correlation length in the strict spectral sense, clearly signals both pseudo-critical temperatures $T_{p_{1}}$ and $T_{p_{2}}$. In general, the spectral correlation length is determined by
@@
whereas pseudo-critical behavior is associated with near-degeneracies between the leading eigenvalues that govern the dominant thermodynamic sectors.
@@
The emergence of two pseudo-critical temperatures in the present model is closely related to the coupling of two diamond units within the same chain. However, increasing the number of diamond units or coupling additional chains does not, in general, guarantee the appearance of further pseudo-critical temperatures. In the coupled twin-diamond chain, the internal structure of the unit cell supports two frustrated manifolds, $\mathrm{FR_{1}}$ and $\mathrm{FR}_{2}$, with different extensive degeneracies, leading to successive entropy-driven free-energy crossings as temperature increases and hence to two well-separated pseudo-transitions. Additional pseudo-transitions may arise only when the enlarged unit cell stabilizes new competing low-energy sectors with distinct residual entropies that survive the coupling, consistently with the criterion discussed in Ref. (bjp20, ).
@@ V Conclusion
The complete phase diagram in the ($J_{2}/J_{1},\mu_{{\rm B}}g_{0}B/J_{1}$) plane shows that the frustrated sectors occupy extended regions between the ordered phases. Because the boundaries between these sectors involve ground-state level crossings accompanied by changes in degeneracy, the associated finite-temperature behavior displays sharp but continuous features. In particular, the entropy and magnetization exhibit steplike but nondiscontinuous variations, while the specific heat and susceptibility develop pronounced finite peaks-typical signatures of pseudo-transition behavior previously reported for other decorated Ising chains.
Altogether, the CTDC provides another exactly solvable example in which competing local configurations and internal degeneracies generate well-defined frustrated regions and clear thermodynamic precursors of pseudo-transitions at low temperature. The analytic structure offers a clear framework for understanding how decoration, sublattice asymmetry, and frustration generate dual pseudo-critical scales in the CTDC model.

### v3

@@ IV.1 Entropy and specific heat
In Fig. 4 is reported the low-temperature entropy landscape in the $(\mu_{{\rm B}}g_{0}B/J_{1},\,k_{{\rm B}}T/J_{1})$ plane for $g_{1}\neq g_{0}$. Panel (left), obtained for $J_{0}/J_{1}=-0.01$, $J_{2}/J_{1}=1.2$ and $g_{1}/g_{0}=0.7$, shows three distinct low-temperature domains that reproduce the structure of the ground-state phase diagram. A narrow green plateau associated with the $\mathrm{FR}_{1}$ phase appears between the FI and $\mathrm{FR}_{2}$ sectors, and its boundary acquires a negative slope due to the asymmetric Zeeman response. In this region the entropy ridges separating the ordered and frustrated sectors mark the emergence of two pseudotransition scales, one related to the competition between FI and $\mathrm{FR}_{1}$ and another associated with the competition between $\mathrm{FR}_{1}$ and $\mathrm{FR}_{2}$. These features become progressively smoother as the temperature increases. Panel (right), corresponding to $J_{0}/J_{1}=-0.01$, $J_{2}/J_{1}=2.4$ and $g_{1}/g_{0}=1.3$, shows the opposite situation where the frustrated boundaries tilt with positive slope, reflecting the opposite imbalance in the magnetic response. The pseudotransition ridges remain visible, indicating that the entropic signatures of the frustrated manifold are robust against variations in the ratio $g_{1}/g_{0}$.
@@
Figure 4: Density plot of the entropy in the $(\mu_{{\rm B}}g_{0}B/J_{1},\,k_{{\rm B}}T/J_{1})$ plane, for a weak ferromagnetic coupling $J_{0}$. (left) $J_{0}/J_{1}=-0.01$, $J_{2}/J_{1}=1.2$ and $g_{1}/g_{0}=0.7$; (right) Similarly for $J_{0}/J_{1}=-0.01$, $J_{2}/J_{1}=2.4$ and $g_{1}/g_{0}=1.3$.
@@
Figure 5 shows the temperature dependence of the entropy and specific heat for two representative parameter sets reported in the caption of Fig.5, illustrating the emergence of two distinct pseudocritical temperatures. These scales are associated with the $\mathrm{FI}$ to $\mathrm{FR}_{1}$ crossover, with pseudocritical temperature given by Eq.(33) at $k_{{\rm B}}T_{p_{1}}/J_{1}\approx 0.0504764$, and with the $\mathrm{FR}_{1}$ to $\mathrm{FR}_{2}$ crossover, with pseudocritical temperature given by Eq. (34) at $k_{{\rm B}}T_{p_{2}}/J_{1}\approx 0.0788948$. Panels (a) and (b) display the entropy curves, which show two smooth but clearly separated increases as temperature rises. The lower increase at $T_{p_{1}}$ corresponds to the partial activation of the $\mathrm{FR}_{1}$ manifold, whereas the higher increase at $T_{p_{2}}$ signals access to the fully frustrated $\mathrm{FR}_{2}$ sector.
@@
Panels (c) and (d) present the corresponding specific heat $C(T)$, which develops two continuous maxima at $T_{p_{1}}$ and $T_{p_{2}}$. The logarithmic insets make the locations of these maxima particularly visible. The first maximum marks the crossover between the ordered FI region and the partially frustrated $\mathrm{FR}_{1}$ configurations, while the second reflects the competition between $\mathrm{FR}_{1}$ and $\mathrm{FR}_{2}$. We can observe that the peak at $T_{p_{1}}$ is higher and sharper than the second peak at $T_{p_{2}}$. This behavior is consistent with the pseudocritical features reported in the literature(sergio18, ; Rojas2019, ; katarina, ; verissimo, ).
@@
The clear separation between these double thermal anomalies shows that the model supports two well defined pseudocritical temperature scales, each associated with a distinct degeneracy change in the ground-state structure.
@@ IV.2 Magnetic properties
Panels (a) and (b) display $\langle s_{k}\rangle$, $\langle S_{k}\rangle$ and $M$ as functions of temperature. All three curves show two smooth but clearly separated steps. The first step, located at the pseudocritical temperature $T_{p_{1}}$, reflects the partial activation of the $\mathrm{FR}_{1}$ configurations, which introduce local frustration while keeping the nodal subsystem partially aligned. The second step, located at $T_{p_{2}}$, marks the activation of the fully frustrated $\mathrm{FR}_{2}$ manifold, where each dimer contributes an additional independent degree of freedom.
@@
Panels (c) and (d) of Fig. 6 show $\chi_{0}$, $\chi_{1}$, and $\chi_{t}$ on a logarithmic scale. All three susceptibilities develop two pronounced peaks at $T_{p_{1}}$ and $T_{p_{2}}$. These peaks coincide with the pseudocritical temperatures identified from the entropy and specific-heat analysis. Although the peaks appear sharp, they remain finite and represent the typical thermal signatures of pseudotransitions in frustrated one-dimensional systems. Again, the anomalies occur at the same temperatures identified from the entropy and specific heat, and the peak at $T_{p_{1}}$ is higher and sharper than the one at $T_{p_{2}}$. This confirms the consistency of the dual pseudotransition behavior in the CTDC model.
@@ IV.3 Eigenvalues properties and correlation length
Figure 7(a) shows the temperature dependence of the relative eigenvalues $|\lambda_{j}|/\lambda_{1}$ for the parameter set $J_{2}/J_{1}=1.24$, $J_{0}/J_{1}=-0.01$, $\mu_{\mathrm{B}}g_{0}B/J_{1}=1.1$, and $g_{1}/g_{0}=0.7$. Although the eigenvalues satisfy the ordering $\lambda_{1}>\lambda_{2}>\lambda_{3}>\lambda_{4}$, the subleading eigenvalue in modulus is not $\lambda_{2}$ but rather $|\lambda_{4}|$ (see appendix A). This behavior is illustrated by the red curve $|\lambda_{4}|/\lambda_{1}$, while the green and blue curves correspond to $\lambda_{2}/\lambda_{1}$ and $\lambda_{3}/\lambda_{1}$, respectively. The ratio $|\lambda_{4}|/\lambda_{1}$ exhibits a broad plateau between the two pseudocritical temperatures, indicating that the same low-energy sector remains dominant throughout this interval. In contrast, the ratio $\lambda_{2}/\lambda_{1}$ displays two pronounced extrema located precisely at the pseudocritical temperatures $T_{p_{1}}$ and $T_{p_{2}}$, reflecting the successive competition between distinct thermodynamic sectors.
For the same parameters, Fig. 7(b) shows the quantities $\xi_{j}=\bigl[\ln(\lambda_{1}/|\lambda_{j}|)\bigr]^{-1}$ as functions of temperature on a logarithmic scale. Since $|\lambda_{4}|$ is the largest subleading eigenvalue in modulus, the spectral correlation length is given by $\xi=\xi_{4}=\bigl[\ln(\lambda_{1}/|\lambda_{4}|)\bigr]^{-1}$. However, this quantity does not exhibit signatures of the pseudocritical behavior. By contrast, the quantity $\xi_{2}=\bigl[\ln(\lambda_{1}/\lambda_{2})\bigr]^{-1}$, although not a correlation length in the strict spectral sense, clearly signals both pseudocritical temperatures $T_{p_{1}}$ and $T_{p_{2}}$. In general, the spectral correlation length is determined by
@@
whereas pseudocritical behavior is associated with near-degeneracies between the leading eigenvalues that govern the dominant thermodynamic sectors.
@@
The emergence of two pseudocritical temperatures in the present model is closely related to the coupling of two diamond units within the same chain. However, increasing the number of diamond units or coupling additional chains does not, in general, guarantee the appearance of further pseudocritical temperatures. In the coupled twin-diamond chain, the internal structure of the unit cell supports two frustrated manifolds, $\mathrm{FR_{1}}$ and $\mathrm{FR}_{2}$, with different extensive degeneracies, leading to successive entropy-driven free-energy crossings as temperature increases and hence to two well-separated pseudotransitions. Additional pseudotransitions may arise only when the enlarged unit cell stabilizes new competing low-energy sectors with distinct residual entropies that survive the coupling, consistently with the criterion discussed in Ref. (bjp20, ).
@@ V Conclusion
The complete phase diagram in the ($J_{2}/J_{1},\mu_{{\rm B}}g_{0}B/J_{1}$) plane shows that the frustrated sectors occupy extended regions between the ordered phases. Because the boundaries between these sectors involve ground-state level crossings accompanied by changes in degeneracy, the associated finite-temperature behavior displays sharp but continuous features. In particular, the entropy and magnetization exhibit steplike but nondiscontinuous variations, while the specific heat and susceptibility develop pronounced finite peaks-typical signatures of pseudotransition behavior previously reported for other decorated Ising chains.
Altogether, the CTDC provides another exactly solvable example in which competing local configurations and internal degeneracies generate well-defined frustrated regions and clear thermodynamic precursors of pseudotransitions at low temperature. The analytic structure offers a clear framework for understanding how decoration, sublattice asymmetry, and frustration generate dual pseudocritical scales in the CTDC model.